OpenStudy (anonymous):

can anyone help pls evaluate 183 - x to 2nd power when x = 5 183-x 2nd power simplify and type and integer or a decimal

OpenStudy (radar):

\[183-x ^{2}\] is that the expression you wish to determine when x = 5?

OpenStudy (radar):

\[183-5^{2}=183-25=158\]

OpenStudy (radar):

or is the expression\[(183-x)^{2}\] (183-5)(183-5) 178 x 178 =31684